OpenStudy (anonymous):

the endpoints of ST are S (-3,2) and T (5,8) find the length of ST how do i do this?

jimthompson5910 (jim_thompson5910):

Use the distance formula to find the distance between S and T

OpenStudy (anonymous):

well i tried doing that and i got =100 then the square root of 10 but i didnt know if it was right or wrong

jimthompson5910 (jim_thompson5910):

Remember the distance formula is \[\Large d = \sqrt{\left(x_{2}-x_{1}\right)^2+\left(y_{2}-y_{1}\right)^2}\]

jimthompson5910 (jim_thompson5910):

ok let me check

jimthompson5910 (jim_thompson5910):

yep I'm getting 10 as well, nice job

jimthompson5910 (jim_thompson5910):

the distance between S and T is 10 units so ST is a segment that is 10 units long

OpenStudy (anonymous):

so its not supposed to be like cordinates? its just the square root of it? i was confused lol

jimthompson5910 (jim_thompson5910):

well you plug the coordinates in to get sqrt(100) = 10

jimthompson5910 (jim_thompson5910):

if that's what you're asking about

OpenStudy (anonymous):

so the answer is just 10 segments?

jimthompson5910 (jim_thompson5910):

10 units yes

OpenStudy (anonymous):

oh okay thanks :)

jimthompson5910 (jim_thompson5910):

np